# Patch: Mitigating PII Leakage in Language Models with Privacy-Aware Targeted Circuit PatcHing

### v1

@@ Abstract
Language models (LMs) may memorize personally identifiable information (PII) from training data, enabling adversaries to extract it during inference. Existing defense mechanisms such as differential privacy (DP) reduce this leakage, but incur large drops in utility. Based on a comprehensive study using circuit discovery to identify the computational circuits responsible PII leakage in LMs, We hypothesize that specific PII leakage circuits in LMs should be responsible for this behavior. Therefore, we propose Patch (Privacy-Aware Targeted Circuit PatcHing), a novel approach that first identifies and subsequently directly edits PII circuits to reduce leakage. Patch achieves better privacy-utility trade-off than existing defenses, e.g., reducing recall of PII leakage from LMs by up to $65\%$. Finally, Patch can be combined with DP to reduce recall of residual leakage of an LM to as low as $0.01\%$. Our analysis shows that PII leakage circuits persist even after the application of existing defense mechanisms. In contrast, Patch can effectively mitigate their impact. [^1] Code and data are publicly available at https://anonymous.4open.science/r/patch-pii-circuits-E7DE
@@ 1 Introduction
Language models (LMs) have demonstrated remarkable advances (Gemma-Team et al., 2024; Grattafiori et al., 2024), yet their tendency to memorize training data poses privacy risks (Kandpal et al., 2022; Buzaglo et al., 2023; Duan et al., 2024; Hayes et al., 2025b). In particular, prior work has shown that LMs can memorize and reproduce personally identifiable information (PII) from their training data (Huang et al., 2022; Kim et al., 2023; Nakka et al., 2024; Borkar et al., 2025). This makes it possible for adversaries with black-box access to a model to expose such information (Lukas et al., 2023).
@@
To address these limitations, we propose Patch (Privacy-Aware Targeted Circuit PatcHing), which allows editing relevant PII circuits to minimize leakage. We first apply mechanistic interpretability to discover the internal computational structures (or “circuits”) in a LM responsible for PII leakage. We then “patch” or edit these circuits to mitigate the leakage. Our contributions are:
@@ 4.3 Base Models
We use several open-weight LMs such as GPT2-Small  (117M), GPT2-Medium  (345M), and GPT2-Large (774M) following prior work (Lukas et al., 2023; Wu et al., 2024), as well as more recent models such as Llama-3.2-1B (Grattafiori et al., 2024), Qwen3-0.6B and Qwen3-1.7B (Team, 2025).
@@ 4.6 Metrics
Figure 2: Comparison of Patch with other defenses: lower perplexity, precision and recall scores are preferred (closer to lower left corner). Patch consistently performs better than prior defenses across all models.
@@ 8 Conclusion
We introduced Patch, a method grounded in mechanistic interpretability for mitigating PII leakage in LMs through targeted circuit patching. Our approach identifies shared computational pathways responsible for PII memorization across PII types and selectively ablates high-importance edges to reduce privacy risks. Empirical results demonstrate that Patch achieves superior privacy-utility trade-offs compared to prior work. Our extensive experiments verify the effectiveness of Patch across multiple model scales and PII categories compared to non-private fine-tuned baselines. These findings demonstrate that targeted circuit interventions can provide effective privacy protection while preserving the model’s core computational pathways.
@@
In this study, we empirically demonstrate that our method substantially improves open-weight LMs in mitigating PII leaks, however we acknowledge our evaluation is limited to a subset of PII types. We hope to extend our proposed method to a broader set of PII in the future. Second, experiments have not been conducted on large models, due to circuit discovery requiring intensive resources and time, however as new methods arise they can plug into our proposed method.

### v2

@@ Abstract
Language models (LMs) may memorize personally identifiable information (PII) from training data, enabling adversaries to extract it during inference. Existing defense mechanisms such as differential privacy (DP) reduce this leakage, but incur large drops in utility. Based on a comprehensive study using circuit discovery to identify the computational circuits responsible PII leakage in LMs, we hypothesize that specific PII leakage circuits in LMs should be responsible for this behavior. Therefore, we propose Patch (Privacy-Aware Targeted Circuit PatcHing), a novel approach that first identifies and subsequently directly edits PII circuits to reduce leakage. Patch achieves better privacy-utility trade-off than existing defenses, e.g., reducing recall of PII leakage from LMs by up to $65\%$. Finally, Patch can be combined with DP to reduce recall of residual leakage of an LM to as low as $0.01\%$. Our analysis shows that PII leakage circuits persist even after the application of existing defense mechanisms. In contrast, Patch can effectively mitigate their impact. [^1] Code and data are publicly available at https://github.com/ssg-research/pii-patch/
@@ 1 Introduction
Language models (LMs) have demonstrated remarkable advances (Gemma-Team and others, 2024; Grattafiori and others, 2024), yet their tendency to memorize training data poses privacy risks (Kandpal et al., 2022; Buzaglo et al., 2023; Duan et al., 2024; Hayes et al., 2025b). In particular, prior work has shown that LMs can memorize and reproduce personally identifiable information (PII) from their training data (Huang et al., 2022; Kim et al., 2023; Nakka et al., 2024; Borkar et al., 2025). This makes it possible for adversaries with black-box access to a model to expose such information (Lukas et al., 2023).
@@
To address these limitations, we propose Privacy-Aware Targeted Circuit PatcHing (Patch), which allows editing relevant PII circuits to minimize leakage. We first apply mechanistic interpretability to discover the internal computational structures (or “circuits”) in a LM responsible for PII leakage. We then “patch” or edit these circuits to mitigate the leakage. Our contributions are:
@@ 4.3 Base Models
We use several open-weight LMs such as GPT2-Small  (117M), GPT2-Medium  (345M), and GPT2-Large (774M) following prior work (Lukas et al., 2023; Wu et al., 2024), as well as more recent models such as Llama-3.2-1B (Grattafiori and others, 2024), Qwen3-0.6B and Qwen3-1.7B (Team, 2025).
@@ 4.6 Metrics
Figure 2: Comparison of Patch with other defenses: lower perplexity, precision and recall scores are preferred (closer to lower left corner).
@@ 8 Conclusion
We introduced Patch, a method grounded in mechanistic interpretability for mitigating PII leakage in LMs through targeted editing. Our approach identifies shared computational pathways responsible for PII memorization across PII types and selectively ablates high-importance edges to reduce privacy risks. Empirical results demonstrate that Patch achieves superior privacy-utility trade-offs compared to prior work. Our extensive experiments verify the effectiveness of Patch across multiple model scales and PII categories compared to non-private fine-tuned baselines. These findings demonstrate that targeted circuit interventions can provide effective privacy protection while preserving the model’s core computational pathways.
@@
In this study, we empirically demonstrate that our method substantially improves open-weight LMs in mitigating PII leaks, however we acknowledge our evaluation is limited to a subset of PII types. We hope to extend our proposed method to a broader set of PII in the future. Finally, experiments have not been conducted on large models, due to circuit discovery requiring intensive resources and time, however as new methods arise they can integrate into our proposed method.